The mineral chalk, which is made up of CaCO3 can be prepared by the reaction of solutions of CaCl2 and Na2CO3


a. Write the balance chemical equation


b. What volume of 0.500 M CaCl2 and 0.250 M Na2CO3 must be mixed to precipitate 8.51 g of CaCO3 ?


1
Expert's answer
2022-05-02T05:32:03-0400

a. CaCl2 + Na2CO3 = CaCO3 + 2NaCl


b. m(CaCl2) = (m(CaCO3) * Mr(CaCl2)) / Mr(CaCO3)

m(CaCl2) = (8.51 g * 110.98 g/mol) / 100.09 g/mole = 9.44 g

Сm = nV = m/MV; V(CaCl2) = m(CaCl2) / C(CaCl2) / M(CaCl2)

V(CaCl2) = 9.44 g / 0.500 M / 110.98 g/mol = 0.17 L


m(Na2CO3) = ((m(CaCO3) * Mr(Na2CO3)) / Mr(CaCO3)

m(Na2CO3) = (8.51 g * 105.99 g/mol) / 100.09 g/mole = 9.01 g

Сm = nV = m/MV; V(Na2CO3) = m(Na2CO3) / C(Na2CO3) / MNa2CO3)

V(Na2CO3) = 9.01 g /0.250 M / 105.99 g/mol = 0.34 L